Night Report 2016-02-23

Observing: Stefan Taubenberger, Tamar Faran, Markus Kromer

Support Team: N. Elias-Rosa

TAT: Joe Lyman, Kate Maguire

Weather conditions

Start of night: clear, humidity ~ 60%, wind~ 2 m/s

Middle of night: clear, humidity ~60%, wind ~ 3 m/s

End of night: clear, humidity ~35%, wind ~ 3 m/s

------------------------------------------------

Calibrations

Evening: PESSTO_SkyFF_eve_grzi, PESSTO_SOFI_specflat, PESSTO_SOFI_FlatJHK, PESSTO_SOFI_SpecArc

Morning: -

------------------------------------------------

00:0 UT: image analysis

00:30: Classification: AT2016ajm, gr13, 900s, seeing ~1.2"

00:50 UT: Follow-up: ASASSN-15og, gr16, 2x2400s, seeing~1.2"

02:35 UT: Follow-up: OGLE15xl, griz, 2700s, seeing ~1.2"

03:35 UT: Standard: L745a, gr11 + gr13 + gr16, 4x100s, seeing ~1.2"

03:55 UT: Classification: Gaia16afj, gr13, 1800s, seeing ~0.8"

04:30 UT: Classification: AT2016aiy,gr13, 900s, seeing ~0.8"

04:55 UT: Switched to SOFI

05:20 UT: Follow-up: SN2016X, BG + RG, 47m, seeing ~0.7"

06:50 UT: Telluric Standard: HIP64944, BG + RG, 9s + 12s, seeing ~0.7"

07:00 UT: Follow-up: SN2016X, JHK, 5m, seeing ~0.7"

07:15 UT: Follow-up: SN2016adj, BG + RG, 18m, 240s + 480s, seeing ~1" *

08:20 UT: Telluric Standard: HIP64550, BG + RG, 9s + 12s, seeing ~0.7"

08:25 UT: Follow-up: SN2016adj, JHK, 10s + 10s + 20s, seeing ~0.7" **

08:45 UT: Standard: EG274, BG + RG, 120s + 240s, seeing ~0.9" ***

06:05 UT: End of the night

* There was a problem with the count down, which didn't stop and went on to negative values (as already happened multiple times in the past month). Due to that the time dedicated to this object took one hour instead of 20-30 minutes.

** There was a problem with the dithering, which forced us to stop and restart. The operators started the OB from the point in which it was affected, and this might cause one or two extra files to appear in the archive. In addition, the H and K counts reach a level 18,000 with the shortest DIT.

*** problem with the acquisition of the flux standard